OpenStudy (anonymous):

Solve for x. 9(x – 2) = 5(2 + x)

OpenStudy (loinducoeur):

Expand the brackets first. \[9x - 18 = 10 +5x\]Then move the 5x to the left side. Remember that the signs change when you move a term to the other side. \[9x -5x -18 = 10\] Then, move the -18 to the right side. The sign also changes.\[9x-5x=10+18\] Now simplify. \[4x = 28\] And finally, divide 28 by 4 to get x on its own. \[x=28\div4=7\]